# Multiplicative and additive systematics in galaxy density fluctuations and clustering measurements

### v1

@@ 3.3 Likelihood Function
[…]
@@ 3.4 Distinguishing between likelihood models
where $\mathcal{L}$ is the likelihood defined in equation (18), $\Theta$ and $\Theta_{0}$ are the unrestricted and restricted parameter spaces, and the hatted variables represent their best fit parameter estimates. So for example, when we want to test whether the additive systematic model is sufficient to describe systematics due to template $t_{i}$, we can define $\Theta_{0}=\{\Theta:b_{i}=0\}$. Similarly when we want to test the sufficiency of a Gaussian likelihood to describe the data, we can define $\Theta_{0}=\{\Theta:\gamma=0\}$. In a likelihood ratio test formulated this way, the null hypothesis is that the data can be adequately described by a simpler model, which usually translates to a model with fewer parameters or restrictions. If the null hypothesis is true, the quantity $\lambda_{\text{LR}}$ defined earlier is $\chi^{2}$ distributed. In order to reject the null hypothesis (thereby preferring a more complex model to describe the data), we need to define some critical value $z$ such that we can reject the null hypothesis if $\lambda_{\text{LR}}>z$. This critical value is often taken to be $z\approx F^{-1}(1-\alpha,r)$, where $F$ is the cumulative distribution function (CDF) of a $\chi^{2}$ distribution, $\alpha$ is the maximum probability of incorrectly rejecting the null hypothesis, and $r$ the number of degrees of freedom. The value of $\alpha$ is selected based on the desired risk of incorrectly rejecting a simpler model for a more complicated one.
@@
[…]
@@ 3.5 Noise Debiasing
where $\text{Var}[\hat{a}_{i}]$ here represents the noise variance of the parameter $a_{i}$. The same considerations and debiasing procedure applies to the $b$ parameters.
Note that failure to debias the squared parameters can lead to significant overcorrection of the two-point function. This is shown quantitatively in Fig. 10 of Appendix B. This illustrates the necessity of removing the noise bias in the estimate of the squared parameters in order to obtain an unbiased estimate of the clustering signal.
@@ 4.1 Mock Extragalactic Galaxy Catalog
[…]
Note that we chose to implement our method directly on the two-point function (in configuration space), which is why we chose to work directly with galaxy positions, rather than looking at contamination in the power spectra. We used a set of 119 publicly-available KiDS-HOD mock catalogs [^3] See available SLICS products at: http://cuillin.roe.ac.uk/~jharno/SLICS/ with total sky coverage of $100\;\text{deg}^{2}$, along with an associated random catalog. The random catalog has a density approximately 10 times higher than that of the mocks. We work with only one low redshift tomographic bin ($0.40.2<z_{\text{spec}}<$ 0.4) as a proof of concept for the methodology presented in this work. Table 1 summarizes key numbers for the mocks used for this work.
@@ 4.2 Mock Systematic Maps
In order to develop and test our method in a realistic scenario, it is important to generate systematics template maps that emulate the structure and distribution of systematics encountered in real data that will bias the observed number density. To do so, we use the Point Spread Function (PSF) area size (a proxy for seeing) as our test case, given that this is a well known observational parameter that can bias the observed galaxy number density depending on how the detection process is carried out, especially for ground-based surveys where regions with worse seeing can exhibit more severe challenges in deblending overlapping galaxy light profiles. We start by measuring the PSF area size autocorrelation function from HSC Public Data Release 1 (Aihara et al., 2018a, b). From a map of seeing size, we can construct a map of the seeing size overdensity, which is what we use to calculate the autocorrelation function. We then fit the correlation function to a declining exponential, as it fits the data quite well in the angular scales between 1 and 500 arcmin, which fall within the scales of interest for measuring clustering. With this fit and the use of CAMB (Lewis et al., 2000), we go from the correlation function to the power spectrum ($w^{tt}\rightarrow C_{\ell}^{tt}$), where $t$ here refers to the seeing size overdensity template we have constructed from the data. We then fit the power spectrum to an exponential function, as it once again provides a good visual fit to the data in the scales we are interested for clustering ($150040<\ell<$ 1500), and find the best fit to be $C_{\ell}^{tt}\propto\exp[-\ell/6]$. However, since the scales probed by HSC are much larger than those we measure in the KiDS-HOD mocks, we adjust the scale of the exponential decay to $1/500$, to emulate the overall behaviour of the correlation function in the range of scales for this analysis. Note that for this validation test, we are not looking to closely fit for the seeing power spectrum, but rather to have an approximate idea of its spatial correlations as a function of scale. We can then use HEALPY (Zonca et al., 2019) to generate full-sky template overdensity maps, and only use the pixels in the same area as our KiDS catalogs.
@@
This figure shows that despite the limited area coverage of the mock catalogs that drove us to use smaller scales than we would in a realistic analysis, the shape of systematic correlations across the scales used is non-trivial and depicts the complexity we can expect for real systematics. Maps can then be generated using a choice from these families, and in practice all maps are normalized to have equal variance. We can see that for a fixed variance, there will be different patterns in the spatial correlations for these families of power spectra. For example, family (iii) has more power for higher $\ell$ modes or smaller spatial scales than most of the other families, while family (v) can be identified as white noise.
@@ 4.3 Measuring Galaxy Clustering and Contamination Parameters
Figure 3: Left: Filled histogram shows the distribution of true galaxy overdensity $\delta_{g}$ for one systematics-free mock catalog from the KiDS-HOD galaxy simulation. The solid lines show the results of MLE fits for the distribution with a skewed Gaussian (green) or Gaussian (purple) likelihood. The dotted lines show the same fit after removing pixels with $\delta_{g}\geq 1.6$ as potential outliers. Right: Normalized histogram of the estimated variance from each fit showed in the left panel. The histogram is constructed from the 250 walkers used in the MCMC. Before removing potential outliers, the estimated variance is notably different between each choice of likelihood. The vertical lines show the measured variance of $\delta_{g}$ for the distribution with (dashed) and without (solid) the $\delta_{g}<1.6$ cut. Removing outliers appears to provide a better visual fit of the true $\delta_{g}$ distribution in the left panel, and the estimated variances for the two likelihood models agree with each other on the right panel. Our two likelihood models differ in their ability to describe the distribution of $\delta_{g}$ values, but removing outliers (to which both are sensitive) leads to agreement between the two models’ estimated variance and the measured variance.
We estimate the contamination parameters described in equation (12) at the level of the one-point function (i.e., the PDF of $\delta_{g}$ values) by sampling the parameter space using a Monte Carlo Markov Chain (MCMC) algorithm with emcee (Foreman-Mackey et al., 2013). The algorithm needs a user-specified number of walkers to explore the parameter space for a certain number of iterations. The number of walkers should generally be at least twice the number of parameters, while the number of iterations needed to converge might depend on the number of pixels used. For this work, we used $250$ walkers and 1500 iterations. The number of iterations and walkers was chosen empirically such that the likelihood function converges within a reasonable error margin. We also confirmed that increasing the number of walkers or iterations did not considerably decrease the integrated autocorrelation time of the MCMC chain.
@@ 4.4 Quantifying Cosmological Impact
We use the full range of scales from the measured clustering (0.06 – 30 arcmin) in this fitting process. Doing this fit for every mock catalog, we can compute the average over all realizations, $\bar{A}$, to obtain a measure of the bias. The covariance matrix $\bm{\mathsf{C}}_{\text{corr}}$ here is taken from all realizations of the corrected clustering signal as detailed in equation (28). We are interested in the estimated bias on the corrected two-point function, $\Delta\bar{A}=\bar{A}-1$, as the success metric to quantify the significance of residual bias in the corrected two-point function. Note that using the estimated $w_{\text{true}}$ and $w_{\text{corr}}$ to estimate $A$ for every mock catalog largely removes the impact of cosmic variance on $A$, reducing the uncertainties in our analysis.
@@ 5 Results
Figure 6: Each panel shows a density plot of the estimated contamination parameters, $\{a_{i},b_{i}\}$, using a Gaussian likelihood, from all 119 mock catalog realizations. The contour regions show the $1\sigma$ and $2\sigma$ regions. The 6 systematics shown (out of a total of 25 contaminants) form a representative sample of the different systematic families (shown by the scaling of their power spectra, $C_{\ell}^{tt}$) used to generate the systematic maps. The vertical and horizontal solid lines show the true parameters used for contamination, with blue (red) representing additive (multiplicative) systematics. The dashed blue line shows the zero point of the $b$ parameter, where an additive systematic should live in parameter space. We see that across all panels the true parameters fall within $1\sigma$ regions of the estimated parameters and we can use the inferred parameters to identify additive and multiplicative systematics.
@@ 5.1 No Systematics
Before investigating systematics mitigation, we first explore the underlying distribution we are trying to model. The likelihood function in equation (18) is based on certain assumptions about the true $\delta_{g}$ distribution. For the case where we cannot assume $|\delta_{g}|\ll 1$, we know that the underlying distribution cannot be Gaussian, as $\delta_{g}$ is mathematically bounded to be greater than $-1$ but can take any positive value. As seen in Fig. 3, we use the true $\delta_{g}$ field from our mock catalogs (filled histogram) and fit its distribution using both a Gaussian and a skewed Gaussian likelihood function. The solid lines in the left-hand panel show that the skewed Gaussian likelihood provides a better fit to the data than the Gaussian likelihood, given the skewed distribution of $\delta_{g}$ values. The right-hand panel shows the difference in the estimated variance between the models. However, we found that this difference was mostly due to potential outliers, which once removed (shown as the dashed lines) provided better fits to the distribution and agreement in the estimated variance. The cut $\delta_{g}<1.6$ was made to test this. Note that in a real cosmological analysis, larger scales will be used and the variance will therefore be smaller.
@@ 5.2 One Systematic
We next test the method for a scenario with a single source of systematic contamination. We consider separately the cases of a single additive and a single multiplicative systematic. The true overdensity map is contaminated using a single systematic template map generated from the power spectrum (i) described in Subsection 4.2 to produce the observed galaxy overdensity map. The observed two-point function is calculated using equation (23). In each realization, we estimate the contamination using the parameters from the MLE approach described in Subsection 3.3, then correct the two-point function using equation (15). We use different versions of the likelihood model in equation (18) to test different assumptions about the distribution of $\delta_{g}$ and its contamination. In Fig. 4, we show the fractional difference between the uncorrected, corrected and true clustering signals to evaluate performance. For context, the systematic correlation function is approximately 2 orders of magnitude smaller than the signal.
As expected, assuming the wrong model for the systematic contamination can result in large residual biases even after applying our systematics mitigation scheme, while choosing the correct model results in removal of the systematic to below the 0.1 percent level. The combined model performs almost identically to choosing the correct model, without substantially increased uncertainties despite the small increase in model complexity (one additional parameter). The similarities in performance between the true and combined model suggest that the combined model is a viable alternative for use in real data. The choice of the likelihood form (Gaussian vs. skewed Gaussian) seems to have no significant impact on performance, so choosing the simpler Gaussian model is preferred in this case.

### v2

@@ 3.3 Likelihood Function
To estimate the contamination parameters, we work directly with galaxy and systematic overdensity maps. We use healpy (Zonca et al., 2019) and HEALPix [^2] http://healpix.sourceforge.net (Górski et al., 2005) to pixelize the data and create such maps from galaxy catalogs. Hence, the notation $\delta_{t,j}$ refers to the average value of $\delta_{t}$ from all galaxies in pixel $j$. The size of the pixels depends on the sample: it should be chosen such that pixels have a non-negligible number of galaxies, but are small enough to retain information about clustering on the smallest scales of interest. In this work we used pixels of NSIDE = $512$, which have an area of $\sim 47$ arcmin2. We use a maximum likelihood estimation approach to estimate the parameters of interest.
@@ 3.4 Distinguishing between likelihood models
where $\mathcal{L}$ is the likelihood defined in equation (18), $\Theta$ and $\Theta_{0}$ are the unrestricted and restricted parameter spaces, and the hatted variables represent their best fit parameter estimates. So for example, when we want to test whether the additive systematic model is sufficient to describe systematics due to template $t_{i}$, we can define $\Theta_{0}=\{\Theta:b_{i}=0\}$. Similarly when we want to test the sufficiency of a Gaussian likelihood to describe the data, we can define $\Theta_{0}=\{\Theta:\gamma=0\}$. In a likelihood ratio test formulated this way, the null hypothesis is that the data can be adequately described by a simpler model, which usually translates to a model with fewer parameters or restrictions. If the null hypothesis is true, the quantity $\lambda_{\text{LR}}$ defined earlier is asymptotically $\chi^{2}$ distributed. In order to reject the null hypothesis (thereby preferring a more complex model to describe the data), we need to define some critical value $z$ such that we can reject the null hypothesis if $\lambda_{\text{LR}}>z$. This critical value is often taken to be $z\approx F^{-1}(1-\alpha,r)$, where $F$ is the cumulative distribution function (CDF) of a $\chi^{2}$ distribution, $\alpha$ is the probability of incorrectly rejecting the null hypothesis, and $r$ the number of degrees of freedom. The value of $\alpha$ is selected based on the desired risk of incorrectly rejecting a simpler model for a more complicated one.
@@
Mocks Information $z$ range $0.40.2<z<$ 0.4 $N_{\text{realizations}}$ 119 $\overline{N}_{\text{gal}}$ 303371 $\overline{n}_{\text{gal}}$ (arcmin-2) 0.84 Number of Random Galaxies 3036000
@@ 3.5 Noise Debiasing
where $\text{Var}[\hat{a}_{i}]$ here represents the noise variance of the parameter $a_{i}$. The same considerations and debiasing procedure applies to the $b$ parameters. In the limit of purely additive systematics, the noise debiasing presented here is in essence equivalent to the debiasing methods developed in mode projection and template subtraction.
Note that failure to debias the squared parameters can lead to significant overcorrection of the two-point function. This is shown quantitatively in Fig. 10 of Appendix B. This illustrates the necessity of removing the noise bias in the estimate of the squared parameters in order to obtain an unbiased estimate of the clustering signal. This particular debiasing method differs from that of methods like ElasticNet, which deal with the bias-variance tradeoff by using data-calibrated L1 and L2 regularization, as mentioned in Sec. 2.3.2. Meanwhile, this approach balances this tradeoff using a direct de-biasing technique similar to that of mode projection and template subtraction.
@@ 4.1 Mock Extragalactic Galaxy Catalog
To validate our methodology we use the KiDS-like Halo Occupation Distribution (HOD) $N$-body simulations (KiDS-HOD hereafter) produced by Harnois-Deraps et al. (2018). This set of galaxy mock catalogs is produced from the wider set of SLICS (Scinet LIght Cone Simulations) described in Harnois-Déraps & van Waerbeke (2015). The 1025 $N$-body simulations used in SLICS are produced by the gravity solver CUBEP3M (Harnois-Déraps et al., 2013), where each run follows $1536^{3}$ particles inside a grid cube of comoving side length $L_{\text{box}}=505\;h^{-1}$ Mpc. The $N$-body simulation produces a full light cone reconstruction, which requires simulating the non-linear evolution of the particles from a desired redshift, as well as generating halo catalogues. The generated halo catalogs are then used, along with a given HOD model, to assign a certain galaxy population to every host halo. The HOD model used to produce these mocks is described in Smith et al. (2017) and it is referred to as the GAMA HOD model in Harnois-Deraps et al. (2018). For the KiDS-HOD galaxy mocks, the HOD model is calibrated to reproduce key properties of the KiDS Survey (Hildebrandt et al., 2016). The galaxies are assigned up to redshift $z_{\text{spec}}=2.0$ and contain photometric redshifts and lensing information as well.
Note that we chose to implement our method directly on the two-point function (in configuration space), which is why we chose to work directly with galaxy positions, rather than looking at contamination in the power spectra. We used a set of 119 publicly-available KiDS-HOD mock catalogs [^3] See available SLICS products at: http://cuillin.roe.ac.uk/~jharno/SLICS/ with total sky coverage of $100\;\text{deg}^{2}$, along with an associated random catalog. The random catalog has a density approximately 10 times higher than that of the mocks. We work with only one low redshift tomographic bin ($0.40.2<z_{\text{spec}}<$ 0.4) as a proof of concept for the methodology presented in this work. Table 1 summarizes key numbers for the mocks used for this work. The catalogs are pixelized with NSIDE = $512$ (pixel area $\sim 49\;\text{arcmin}^{2}$), meaning that a typical pixel has $\sim 40$ galaxies.
@@ 4.2 Mock Systematic Maps
In order to develop and test our method in a realistic scenario, it is important to generate systematics template maps that emulate the structure and distribution of systematics encountered in real data that will bias the observed number density. To do so, we use the Point Spread Function (PSF) area size (a proxy for seeing) as our test case, given that this is a well known observational parameter that can bias the observed galaxy number density depending on how the detection process is carried out, especially for ground-based surveys where regions with worse seeing can exhibit more severe challenges in deblending overlapping galaxy light profiles. We start by measuring the PSF area size autocorrelation function from HSC Public Data Release 1 (Aihara et al., 2018a, b). From a pixelized map of seeing size, with NSIDE = 512, we can construct a map of the seeing size overdensity, which is what we use to calculate the autocorrelation function using the estimator described in Sec. 4.3. We then fit the correlation function to a declining exponential, as it fits the data quite well in the angular scales between 1 and 500 arcmin, which fall within the scales of interest for measuring clustering. With this fit and the use of CAMB (Lewis et al., 2000), we go from the correlation function to the power spectrum ($w^{tt}\rightarrow C_{\ell}^{tt}$), where $t$ here refers to the seeing size overdensity template we have constructed from the data. We then fit the power spectrum to an exponential function, as it once again provides a good visual fit to the data in the scales we are interested for clustering ($150040<\ell<$ 1500), and find the best fit to be $C_{\ell}^{tt}\propto\exp[-\ell/6]$. However, since the scales probed by HSC are much larger than those we measure in the KiDS-HOD mocks, we adjust the scale of the exponential decay to $1/500$, to emulate the overall behaviour of the correlation function in the range of scales for this analysis. Note that for this validation test, we are not looking to closely fit for the seeing power spectrum, but rather to have an approximate idea of its spatial correlations as a function of scale. We can then use HEALPY (Zonca et al., 2019) to generate full-sky template overdensity maps with NSIDE = 512, and only use the pixels in the same area as our KiDS catalogs.
@@
This figure shows that despite the limited area coverage of the mock catalogs that drove us to use smaller scales than we would in a realistic analysis, the shape of systematic correlations across the scales used is non-trivial and depicts the complexity we can expect for real systematics. Maps can then be generated using a choice from these families, and in practice all maps are normalized to have equal variance. We can see that for a fixed variance, there will be different patterns in the spatial correlations for these families of power spectra. For example, family (iii) has less power for higher $\ell$ modes or smaller spatial scales than most of the other families, while family (v) can be identified as white noise.
@@ 4.3 Measuring Galaxy Clustering and Contamination Parameters
Figure 3: Filled histogram shows the distribution of true galaxy overdensity $\delta_{g}$ for one systematics-free mock catalog from the KiDS-HOD galaxy simulation. The solid lines show the results of maximum likelihood estimation fits for the distribution with a skewed Gaussian (green) or Gaussian (purple) likelihood. All contamination parameters are fixed to zero while fitting, while $\sigma$ is free to vary in both likelihood forms. The parameter $\gamma=0$ for the Gaussian likelihood but is free to vary for the skewed Gaussian. The dotted lines show the same fit after removing pixels with $\delta_{g}\geq 1.6$ as potential outliers.
We estimate the contamination parameters described in equation (12) at the level of the one-point function (i.e., the PDF of $\delta_{g}$ values) by sampling the parameter space using a Monte Carlo Markov Chain (MCMC) algorithm with emcee (Foreman-Mackey et al., 2013) and choosing the parameters corresponding to the maximum likelihood point. Even though parameters can be more efficiently obtained by using an optimizer (such as gradient descent), we use MCMC to obtain estimates of the parameter noise from single mock realizations (without the need to assume a Gaussian likelihood) and compare them to those from multiple mock realizations. The algorithm needs a user-specified number of walkers to explore the parameter space for a certain number of iterations. The number of walkers should generally be at least twice the number of parameters, while the number of iterations needed to converge might depend on the number of pixels used. For this work, we used $250$ walkers and 1500 iterations. We evaluate convergence heuristically by running the MCMC for 5000 iterations, on a single mock realization, using the maximum number of parameters considered in this work. We confirm that the mean and variance of each parameter, calculated across all chains, stabilizes to the 5% level on average after 1500 iterations. We also confirmed that increasing the number of walkers or iterations did not considerably decrease the integrated autocorrelation time of the MCMC chain, which was approximately 90 for a 52 parameter inference.
@@ 4.4 Quantifying Cosmological Impact
We use the full range of scales from the measured clustering (0.06 – 30 arcmin) in this fitting process. Doing this fit for every mock catalog, we can compute the average over all realizations, $\bar{A}$, to obtain a measure of the bias. The covariance matrix $\bm{\mathsf{C}}_{\text{corr}}$ here is taken from all realizations of the corrected clustering signal and is calculated using equation (24). We are interested in the estimated bias on the corrected two-point function, $\Delta\bar{A}=\bar{A}-1$, as the success metric to quantify the significance of residual bias in the corrected two-point function. Note that using the realization-specific $w_{\text{true}}$ and $w_{\text{corr}}$ to estimate $A$ for every mock catalog largely removes the impact of cosmic variance on $A$, reducing the uncertainties in our analysis.
4.5 Summary: Step-by-Step Procedure
We summarize the previous subsections with a step-by-step procedure for generating and analyzing the data of a single mock catalog.
  1. 1.
    Measure the true correlation function of the extragalactic galaxy catalog described in Sec. 4.1 and generate pixelized maps with $\text{NSIDE}=512$.
  2. 2.
    Generate $N_{\text{sys}}$ full-sky template maps using HEALPY from the desired selection of the 5 systematics families described in Sec. 4.2. Mask the full-sky map to only cover the same area as the galaxy mock catalog. Compute the correlation function for all systematics by assigning each galaxy the corresponding template values for the pixel in which the galaxy is located.
  3. 3.
    From the pixelized versions of the galaxy field ($\delta_{g}$) and systematics maps ($\delta_{t}$), we generate the contaminated galaxy field ($\hat{\delta}_{g}$) using equation  (12) from a chosen set of contamination parameters. We then calculate the contaminated clustering signal using equation  (14).
  4. 4.
    With the observed galaxy overdensity field and systematics maps as inputs, the contamination parameters are jointly estimated as prescribed in Sec. 4.3 by using a MCMC algorithm with the chosen likelihood model (Gaussian or skewed Gaussian) and contamination model (additive, multiplicative, or combined).
  5. 5.
    The contaminated clustering signal is corrected using equation (15).
This procedure is followed for all 119 mock realizations, yielding an ensemble of results used to quantify decontamination quality, parameter estimation and noise, and cosmological impact.
@@ 5 Results
Figure 6: Each panel shows a density plot of the estimated contamination parameters, $\{a_{i},b_{i}\}$, using a Gaussian likelihood, from all 119 mock catalog realizations. In other words, every point on a particular panel represents the maximum likelihood estimate from all MCMC samples of a single mock catalog. The contour regions show the $1\sigma$ (68%) and $2\sigma$ (95%) regions. The 6 systematics shown (out of a total of 25 contaminants) form a representative sample of the different systematic families (shown by the scaling of their power spectra, $C_{\ell}^{tt}$) used to generate the systematic maps. The vertical and horizontal solid lines show the true parameters used for contamination, with blue (red) representing additive (multiplicative) systematics. The dashed blue line shows the zero point of the $b$ parameter, where an additive systematic should live in parameter space. We see that across all panels the true parameters fall within $1\sigma$ regions of the estimated parameters and we can use the inferred parameters to identify additive and multiplicative systematics.
@@ 5.1 No Systematics
Before investigating systematics mitigation, we first explore the underlying distribution we are trying to model. The likelihood function in equation (18) is based on certain assumptions about the true $\delta_{g}$ distribution. For the case where we cannot assume $|\delta_{g}|\ll 1$, we know that the underlying distribution cannot be Gaussian, as $\delta_{g}$ is mathematically bounded to be greater than $-1$ but can take any positive value. As seen in Fig. 3, we use the true $\delta_{g}$ field from our mock catalogs (filled histogram) and fit its distribution using both a Gaussian and a skewed Gaussian likelihood function. For both likelihoods we fix any contamination parameters {$a,b$} to zero and allow $\sigma$ to vary, while we fix $\gamma=0$ for the Gaussian likelihood and let it vary for the skewed Gaussian. The solid lines in the left-hand panel show that the skewed Gaussian likelihood provides a better fit to the data than the Gaussian likelihood, given the skewed distribution of $\delta_{g}$ values.
However, we found that this difference was mostly due to potential outliers, which once removed (shown as the dashed lines) provided better fits to the distribution and agreement in the estimated variance. Physically the outliers represent regions of very high galaxy density. The cut $\delta_{g}<1.6$ was made to test this. Note that in a real cosmological analysis, larger scales will be used and the variance will therefore be smaller.
@@ 5.2 One Systematic
We next test the method for a scenario with a single source of systematic contamination. We consider separately the cases of a single additive and a single multiplicative systematic. The true overdensity map is contaminated using a single systematic template map generated from the power spectrum (i) described in Subsection 4.2 to produce the observed galaxy overdensity map. The observed two-point function is calculated using equation (23). In each realization, we estimate the contamination using the parameters from the maximum likelihood estimation approach described in Subsection 3.3 by fitting for a single systematic map, then correct the two-point function using equation (15). We use different versions of the likelihood model in equation (18) to test different assumptions about the distribution of $\delta_{g}$ and its contamination. In Fig. 4, we show the fractional difference between the uncorrected, corrected and true clustering signals to evaluate performance. For context, the systematic correlation function is approximately 2 orders of magnitude smaller than the signal.
As expected, assuming the wrong model for the systematic contamination can result in large residual biases even after applying our systematics mitigation scheme, while choosing the correct model results in removal of the systematic to below the 0.1 percent level. The combined model performs almost identically to choosing the correct model, without substantially increased uncertainties despite the small increase in model complexity (one additional parameter). The similarities in performance between the true and combined model suggest that the combined model is a viable alternative for use in real data. The choice of the likelihood form (Gaussian vs. skewed Gaussian) seems to have no significant impact on performance, so choosing the simpler Gaussian model is preferred in this case. However, given the generalized contamination model chosen, we felt it was worth revisiting the question of likelihood choice and test its impact directly. The marginal differences between likelihoods found here reinforces previous works’ conclusions regarding the Gaussian approximation for systematics mitigation in galaxy clustering (Weaverdyck & Huterer, 2021; Rezaie et al., 2021).